OpenStudy (anonymous):

Let R , the real line be endowned with the discrete topology. Which of the following subsets of R is dense in R

OpenStudy (anonymous):

Q Ritself Qc All singletons

OpenStudy (steve816):

Wow, this is some advanced stuff. Just wondering, what branch of mathematics is this?

OpenStudy (anonymous):

topology

OpenStudy (steve816):

Nice, wish there was topology class in high school, did you learn about the banach-tarski paradox?

OpenStudy (anonymous):

no

OpenStudy (steve816):

Well, it's pretty interesting. I don't think anyone in openstudy knows anything about topology though.

OpenStudy (anonymous):

someone does

OpenStudy (steve816):

At least not at 3 AM in the morning lol

OpenStudy (anonymous):

yes

OpenStudy (steve816):

So... how hard is topology?

OpenStudy (anonymous):

in a way very

OpenStudy (zzr0ck3r):

All sets are open in the discrete topology. Consider any any set \(U\ne X\). Then \(U^C\) is an open set disjoint from \(U\). This shows that the only dense set is the whole space.

OpenStudy (zzr0ck3r):

\(U\) is dense means that for any \(x\) in the space \(X=\mathbb{R} _D\), either \(x\) is in \(U\) or \(x\) is a limit point of \(U\). So we take an element \(x\) that is not in \(U\) and see if it is a limit point of \(U\). If \(x\) is a limit point of \(U\) then every open neighborhood around \(x\) intersects \(U\). But since every set is open, due to the topology, we have that \(U^C\) itself is open and contains \(x\) (remember \(x\notin U\)) and of course \(U\cap U^C=\emptyset\). So \(U\) contains all of its limit points. Since \(U\) was not all of \(\mathbb{R}_D\) we have that the only dense set is the entire space.

OpenStudy (zzr0ck3r):

Clearly the entire space is dense because it contains all of its own points!

OpenStudy (zzr0ck3r):

@steve816 it really is not, and it is a fun new (new to me...)way of looking at things :)

OpenStudy (anonymous):

so, which option is correct

OpenStudy (anonymous):

is it R?

OpenStudy (zzr0ck3r):

It should be very clear. Read it and tell me if you don't understand.

OpenStudy (zzr0ck3r):

I want you to at least be sure yourself. I did not write all of that out so that you could ask which option it is, click a button, and move on.

OpenStudy (anonymous):

ok, from what you said, i its R

OpenStudy (anonymous):

@zzr0ck3r